Determine the displacement vector that must be added to the displacement(25i - 16j) metre to give a displacement of 7.0 metres pointing in the x+ direction

1
Expert's answer
2022-05-02T13:43:30-0400

"|\vec r|=\sqrt{25^2+16^2}=29.7,"

"x=29.7-7.0=22.7~m."